# Deriving Causal Order from Single-Variable Interventions: Guarantees & Algorithm

### v1

@@ 9 Empirical results
Figure 2: Comparison of the performance of the baselines and of our model Intersort across diverse data domains as presented (linear, RFF, NN and GRN data), for $30$ variables. The x-axis corresponds to the fraction of variables that have been targeted by an intervention. The y-axis is the performance of causal ordering prediction as measured by the $D_{top}$ metric (see section 3, lower is better). The violins are order from left to right: EASE, PC, GIES, Intersort. Results for $10$ variables can be found in the appendix (fig. 3).
@@
We now evaluate Intersort on simulated empirical data and compare its performance to various baselines. The graphs are simulated from a Erdos-Renyi distribution (Erdős et al., 1960), with an expected number of edges per variable $c\in\{1,2\}$. We follow a setup close to (Lorch et al., 2022) to simulate data from linear and random Fourier features (RFF) additive functional relationships. We also apply the models to simulated single-cell data from the SERGIO (Dibaeinia & Sinha, 2020) model, using the code from Lorch et al. (2022) (MIT License, v1.0.5). Lastly, we apply on neural network functional data following the setup of Brouillard et al. (2020) and using the implementation of Nazaret et al. (2023) (MIT License, v0.1.0). We run the models on varying levels of intervention, where the ratio of intervened variables is in $\{0\%,25\%,50\%,75\%,100\%\}$. The data is always standardized based on the mean and variance of the observational dataset, which removes the Varsortability (Reisach et al., 2021) artefact in the data. For the RFF and linear domain, the distribution of the noise is chosen uniformly at random from uniform Gaussian (noise scale independent from the parents), heteroscedastic Gaussian (noise scale functionally dependent on the parents), and Laplace. For the neural network domain, the noise distribution is Gaussian with a fixed variance. We run on $10$ simulated datasets for each domain and each ratio of intervened variables. We simulate $5000$ samples for the observational datasets and $100$ samples for each interventions, which is a setting similar to real single cell transcriptomics datasets (Replogle et al., 2022).
@@
Choosing appropriate baselines to compare to is not trivial, given that, to our knowledge, our setting of predicting the causal order from interventional data has not been considered in the literature, as the focus in prior literature has been on observational data (Reisach et al., 2021; Bühlmann et al., 2014; Rolland et al., 2022; Shimizu et al., 2011). As such, we construct causal orders from the CPDAG predicted by two standard methods, namely PC (Spirtes et al., 2000) and GIES (Hauser & Bühlmann, 2012), by creating a DAG from the oriented edges and then computing the topological order of this DAG. PC is a constrained based method relying on conditional independence tests to estimate the structure of the graph. We run the implementation of Zheng et al. (2024) (MIT License, causal-learn v0.1.3). GIES is an extension to interventional data of the GES model (Chickering, 2002), which is a score based method that greedily adds and removes edges to the estimated CPDAG. We use the Gaussian BIC score, and run the package implementation of Gamella (2022) (BSD 3-Clause License, v0.0.1). Additionally, we compare to EASE (Gnecco et al., 2021), which is a method that aims to learn the causal order from observational data, leveraging the insight that extreme tail-values reveal the causal order. Given that interventions can lead to extreme values not present in the observational data, we also compare to this method. For EASE, we run our own Python implementation of the algorithm. For Intersort, we use $\epsilon=0.3$ for the linear, RFF and neural network domain, and $\epsilon=0.5$ for the single-cell domain. An open-source implementation of Intersort is available at https://github.com/MathieuChevalley/Intersort (Apache 2.0 License). We compute the Wassertein distances with the SciPy python package (Virtanen et al., 2020).
@@
The results for each method on the different types of data considered are displayed in fig. 2, where the distribution of the $D_{top}$ score of each model is plotted against the ratios of intervened variables considered. As can be observed, Intersort performs well on all data domains, and shows decreasing error as more interventions are available, exhibiting the model’s capability to capitalize on the interventional information to recover the causal order across diverse settings. Furthermore, its performance at $50\%$ of intervened variables is close to its performance at higher fractions of intervened variables. Compared to the baselines, only GIES in the linear domain performs better. These experiments demonstrate that $\epsilon$-interventional faithfulness is fulfilled by a diverse set of data types, and that this property can be robustly exploited to recover causal information. Only when strong assumptions about the data distributions are fulfilled, such as linear data for GIES, can better results be obtained.
@@ 10 Discussion and conclusion
We note that Intersort’s performance could still be improved in various ways. For example, a more scalable or closer to the optimum approximation algorithm could be designed. Also, the optimal parameter $\epsilon$ could be chosen in a data dependent way instead of using a fixed value. Lastly, other statistical distances, especially ones that converge faster to 00 in terms of number of samples when the two distributions are equal, could lead to better results. We leave such potential improvements as future work. Another question that may need further investigation is the $\epsilon$-interventional faithfulness assumption. We argue that it is not any stronger than common assumption in the literature such as faithfulness, and lemma 3 shows that it already covers a sizeable set of SCMs. Furthermore, our empirical results demonstrate that it probably holds for a rather large class of SCMs. We also derive theoretical results for a relaxation of this assumption in section 6. An interesting direction of theoretical work would be to further analyse the intricate relationship between the statistical distance $D$, the number of variables $d$, the parameter $\epsilon$, the interventional random variables $\tilde{N}$, and the sample size for non-asymptotic settings. For example, out framework makes it amenable to study the strength of the interventions needed to reveal the causal descendants above a statistically significant threshold.
@@
In this study, we introduced Intersort, an innovative algorithm designed to uncover the causal order from interventional data by optimizing for a new proposed score on causal orders. We introduced the $\epsilon$-interventional faithfulness assumption and proved that interventional datasets fulfilling it have strong guarantees in terms of upper-bounds on the expected error of the optimum of our score. The performance of Intersort and the validity of our theoretical findings was extensively demonstrated on a diverse set of simulated datasets, across functional relationships, noise types and graph densities. Intersort has demonstrated superior performance when compared to established benchmarks such as PC, GIES, and EASE. The robustness and versatility of Intersort underscore the potential of the $\epsilon$-interventional faithfulness assumption to reshape causal inference methodologies. We envision that this work should spearhead new model development based on our proposed perspective on dataset with high numbers of single variable interventions. This includes downstream tasks such as causal discovery, for example by reconstructing the graph given the estimated causal order. It could inform active intervention selection as well, as designing a policy to choose which variables to target next may lead to better guarantees than our results based on interventions selected uniformly at random. All these directions and method developments can have real world impact on critical application domains such as biology, from which this work and its assumptions were inspired.
Acknowledgements and disclosure
@@ Appendix A Proofs of Theorems
Let $\mathbf{D}$ be a $d\times d$ matrix containing all the distances, where $\mathbf{D}(i,j)=D\left(P_{X_{j}}^{\mathcal{C},(\emptyset)},P_{X_{j}}^{\mathcal{C},do(X_{i}:=\tilde{N}_{i})}\right)$.
@@ Appendix G Computational resources
The simulations of fig. 1 were run on an Apple M1 Macbook pro. The experiments for figs. 2 and 3 were run on a cluster with 20 CPUs, 16Gb of memory per CPU.

### v3

@@ 10 Empirical results
Figure 2: Comparison of the performance of the baselines and of our model Intersort across diverse data domains as presented (linear, RFF, NN and GRN data), for $30$ variables. The x-axis corresponds to the fraction of variables that have been targeted by an intervention. The y-axis is the performance of causal ordering prediction as measured by the $D_{top}$ metric (see section 3, lower is better). The violins are order from left to right: EASE, PC, DCDI, GIES, Intersort. Results for $10$ and $100$ variables can be found in the appendix (figs. 7 and 8).
@@
We now evaluate Intersort on simulated empirical data and compare its performance to various baselines. The graphs are simulated from a Erdos-Renyi distribution (Erdős et al., 1960), with an expected number of edges per variable $c\in\{1,2\}$. We follow a setup close to (Lorch et al., 2022) to simulate data from linear and random Fourier features (RFF) additive functional relationships. We also apply the models to simulated single-cell data from the SERGIO (Dibaeinia & Sinha, 2020) model, using the code from Lorch et al. (2022) (MIT License, v1.0.5). Lastly, we apply on neural network functional data following the setup of Brouillard et al. (2020) and using the implementation of Nazaret et al. (2023) (MIT License, v0.1.0). We run the models on varying levels of intervention, where the ratio of intervened variables is in $\{25\%,50\%,75\%,100\%\}$. The data is always standardized based on the mean and variance of the observational dataset, which removes the Varsortability (Reisach et al., 2021) artefact in the data. For the RFF and linear domain, the distribution of the noise is chosen uniformly at random from uniform Gaussian (noise scale independent from the parents), heteroscedastic Gaussian (noise scale functionally dependent on the parents), and Laplace. For the neural network domain, the noise distribution is Gaussian with a fixed variance. We run on $10$ simulated datasets for each domain and each ratio of intervened variables. We simulate $5000$ samples for the observational datasets and $100$ samples for each interventions, which is a setting similar to real single cell transcriptomics datasets (Replogle et al., 2022).
@@
Choosing appropriate baselines to compare to is not trivial, given that, to our knowledge, our setting of predicting the causal order from interventional data has not been considered in the literature, as the focus in prior literature has been on observational data (Reisach et al., 2021; Bühlmann et al., 2014; Rolland et al., 2022; Shimizu et al., 2011). As such, we construct causal orders from the CPDAG predicted by two standard methods, namely PC (Spirtes et al., 2000) and GIES (Hauser & Bühlmann, 2012), by creating a DAG from the oriented edges and then computing the topological order of this DAG. PC is a constrained based method relying on conditional independence tests to estimate the structure of the graph. We run the implementation of Zheng et al. (2024) (MIT License, causal-learn v0.1.3). GIES is an extension to interventional data of the GES model (Chickering, 2002), which is a score based method that greedily adds and removes edges to the estimated CPDAG. We use the Gaussian BIC score, and run the package implementation of Gamella (2022) (BSD 3-Clause License, v0.0.1). Additionally, we compare to EASE (Gnecco et al., 2021), which is a method that aims to learn the causal order from observational data, leveraging the insight that extreme tail-values reveal the causal order. Given that interventions can lead to extreme values not present in the observational data, we also compare to this method. For EASE, we run our own Python implementation of the algorithm. Lastly, we compare to DCDI (Brouillard et al., 2020), which is a continuously differentiable causal discovery model that can leverage interventional data. For Intersort, we use $\epsilon=0.3$ for the linear, RFF and neural network domain, and $\epsilon=0.5$ for the single-cell domain. We set $c=0.5$. We compute the Wassertein distances with the SciPy python package (Virtanen et al., 2020).
@@
The results for each method on the different types of data considered are displayed in fig. 2, where the distribution of the $D_{top}$ score of each model is plotted against the ratios of intervened variables considered. As can be observed, Intersort performs well on all data domains, and shows decreasing error as more interventions are available, exhibiting the model’s capability to capitalize on the interventional information to recover the causal order across diverse settings. Compared to the baselines, only GIES in the linear domain and DCDI in the neural network domain at $25\%$ and $50\%$ perform better. These experiments demonstrate that $\epsilon$-interventional faithfulness is fulfilled by a diverse set of data types, and that this property can be robustly exploited to recover causal information. Only when strong assumptions about the data distributions are fulfilled, such as linear data for GIES, can better results be obtained in low intervention settings.
@@ 11 Discussion and conclusion
We note that Intersort’s performance could still be improved in various ways. For example, a more scalable or closer to the optimum approximation algorithm could be designed. Scalability of Intersort in its current form is mainly constrained by the LOCALSEARCH component, for which alternative optimization techniques could be employed to enhance efficiency and manage larger variable sets. Also, the optimal parameter $\epsilon$ could be chosen in a data dependent way instead of using a fixed value. Lastly, other statistical distances, especially ones that converge faster to 00 in terms of number of samples when the two distributions are equal, could lead to better results. We leave such potential improvements as future work. Another question that may need further investigation is the $\epsilon$-interventional faithfulness assumption. We argue that it is not any stronger than common assumption in the literature such as faithfulness, and lemma 3 shows that it already covers a sizeable set of SCMs. Furthermore, our empirical results demonstrate that it probably holds for a rather large class of SCMs. We also derive theoretical results for a relaxation of this assumption in section 7. An interesting direction of theoretical work would be to further analyze the intricate relationship between the statistical distance $D$, the number of variables $d$, the parameter $\epsilon$, the interventional random variables $\tilde{N}$, and the sample size for non-asymptotic settings. For example, our framework makes it amenable to study the strength of the interventions needed to reveal the causal descendants above a statistically significant threshold.
@@
In this study, we introduced Intersort, an innovative algorithm designed to uncover the causal order from interventional data by optimizing for a new proposed score on causal orders. We introduced the $\epsilon$-interventional faithfulness assumption and proved that interventional datasets fulfilling it have strong guarantees in terms of upper-bounds on the expected error of the optimum of our score. The performance of Intersort and the validity of our theoretical findings was extensively demonstrated on a diverse set of simulated datasets, across functional relationships, noise types and graph densities. Intersort has demonstrated superior performance when compared to established benchmarks such as PC, GIES, DCDI and EASE. The robustness and versatility of Intersort underscore the potential of the $\epsilon$-interventional faithfulness assumption to reshape causal inference methodologies. We envision that this work should spearhead new model development based on our proposed perspective on dataset with high numbers of single variable interventions. This includes downstream tasks such as causal discovery, for example by reconstructing the graph given the estimated causal order. It could inform active intervention selection as well, as designing a policy to choose which variables to target next may lead to better guarantees than our results based on interventions selected uniformly at random. All these directions and method developments can have real world impact on critical application domains such as biology, from which this work and its assumptions were inspired.
Acknowledgments and disclosure
@@ Appendix A Proofs of Theorems
Let $\mathbf{D}$ be a $d\times d$ matrix containing all the distances, where $\mathbf{D}(i,j)=D\left(P_{X_{j}}^{\mathcal{C},(\emptyset)},P_{X_{j}}^{\mathcal{C},do(X_{i}:=\tilde{N}_{i})}\right)$. For the case of every variables being intervened on, we can set $c=0$.
@@ Appendix C Pseudocodes
Algorithm 1 Complete algorithm
1:procedure Intersort( $\mathcal{S}$: score function eq. 1, $\epsilon$,
$D$: statistical distance function,
$\mathcal{I}$: index set of intervened variables,
$P_{X}^{\mathcal{C},(\emptyset)}$: observational distribution,
$\mathcal{P}_{int}$: interventional distributions )
2:     Initialize $\mathbf{D}$ as a zero matrix of size $d\times d$
3:     for each $(i,j)\in[1..d]$ do
4:         if $i\in\mathcal{I}$ then
6:         end if
7:     end for
8:     $\pi_{init}\leftarrow\textsc{sortranking}(\mathbf{D},\epsilon)$
10:     return $\pi_{opt}$
11:end procedure
@@ Appendix F Additional experiments
F.1 Simulations
Additional results for the simulation evaluations. Figure 3 shows the performance of Intersort for scale-free networks of size $5$ and $30$. Figures 4 and 5 show the performance of SORTRANKING on large scale settings of $1000$ and $2000$ variables, for both scale-free and Erdos-Renyi networks. Figure 6 presents a possible heuristic for selecting variables to intervene on, and comparing to a random selection.
(a) Simulation with 5 variables
(b) Simulation with 30 variables
Figure 3: Simulation and comparison between the bounds of theorem 4 for scale-free networks, and between Intersort and the exact $\pi_{opt}$. For each setting, we draw $20$ graphs per setting following a Barabasi-Albert scale-free distribution, with average edge per variable in $\{1,2,3\}$. A setting is the tuple $(p_{int},p_{e})$, where $p_{e}=\frac{2\mathrm{E}(\#edges)}{d(d-1)}$. Then, for each graph, we run the algorithm on $10$ configurations, where each configuration corresponds to a draw of the targeted variables following $p_{int}$. For both $5$ and $30$ variables, we have $p_{int}\in\{0.25,0.33,0.5,0.66,0.75\}$. The settings are ordered on the x-axis following what we call the effective intervention ratio $\frac{p_{int}}{\sqrt{p_{e}}}$. We observe that the error is approximately monotonic when ordered by the effective intervention ratio.
(a) Simulation with 1000 variables
(b) Simulation with 2000 variables
Figure 4: Simulation and comparison between the bounds of theorem 4 for scale-free networks and SORTRANKING. For each setting, we draw $2$ graphs per setting following a Barabasi-Albert scale-free distribution, with average edge per variable in $\{1,2,3\}$. A setting is the tuple $(p_{int},p_{e})$, where $p_{e}=\frac{2\mathrm{E}(\#edges)}{d(d-1)}$. Then, for each graph, we run the algorithm on $1$ configuration, where each configuration corresponds to a draw of the targeted variables following $p_{int}$. For both $5$ and $30$ variables, we have $p_{int}\in\{0.25,0.33,0.5,0.66,0.75\}$. The settings are ordered on the x-axis following what we call the effective intervention ratio $\frac{p_{int}}{\sqrt{p_{e}}}$. Even though the objective can be approximately solved at scale, we observe room for improvement as the performance of SORTRANKING is above the upper-bound in many settings.
(a) Simulation with 1000 variables
(b) Simulation with 2000 variables
Figure 5: Simulation and comparison between the two bounds and SORTRANKING. For each setting, we draw $2$ graphs per setting, where a setting is the tuple $(p_{int},p_{e})$. Then, for each graph, we run the algorithm on $1$ configurations, where each configuration corresponds to a draw of the targeted variables following $p_{int}$. For both $5$ and $30$ variables, we have $p_{int}\in\{0.25,0.33,0.5,0.66,0.75\}$. In the $1000$ variables setting on the left, we have $p_{e}\in\{0.005,0.002,0.001\}$. In the $20000$ variables settings on the right, we have $p_{e}\in\{0.0001,0.00005,0.00002\}$. The settings are ordered on the x-axis following what we call the effective intervention ratio $\frac{p_{int}}{\sqrt{p_{e}}}$. Even though the objective can be approximately solved at scale, we observe room for improvement as the performance of SORTRANKING is above the upper-bound in many settings.
Figure 6: Simulation of scale free networks with average edge per variable in $\{1,2,3\}$ for 30 variables. We report the performance of Intersort with various intervention selection policies. We evaluate three policies: selecting the 10 variables with the most children, selecting 10 variables at random, and selecting the 10 variables with the fewest children. This corresponds to a setting with $30\%$ of intervened variables. We can thus hypothesize that a potential approach for intervention selection is to estimate the connectivity of the non-intervened variables.
F.2 Empirical data
Results for the empirical evaluation for the smaller setting of $10$ variables (fig. 7). Figure 8 shows the performance of SORTRANKING for $100$ variables. Figure 9 analyzes the empirical difference in training time for the baselines on the neural network data type.
@@ F.2 Empirical data
(a) Linear 100 variables
(b) RFF 100 variables
(c) NN 100 variables
(d) GRN 100 variables
Figure 8: Performance of our model SORTRANKING across diverse data domains as presented (linear, RFF, NN and GRN data), for $100$ variables. The x-axis corresponds to the fraction of variables that have been targeted by an intervention. The y-axis is the performance of causal ordering prediction as measured by the $D_{top}$ metric (see section 3, lower is better).
(a) Training Time for $d=10$
(b) Training Time for $d=30$
(c) Relative Change ($d=30$ vs $d=10$)
Figure 9: Training time analysis across different configurations for the NN data type. Plots (a) and (b) show the training time on a logarithmic scale for $d=10$ and $d=30$, respectively. Plot (c) shows the relative change in training time between $d=10$ and $d=30$, also on a logarithmic scale.
F.3 Qualitative example on real-world data
PKCPKAP38JnkMekRafErkPIP2PlcgAktPIP3
Figure 10: Output of Intersort on the flow cytometry data set of the Sachs et al. (2005) dataset, which measures the level of expression of phosphoproteins and phospholipids in human cells. Out of the $11$ proteins, $5$ (in red) where intervened on using reagents to activate or inhibit the measured proteins. The dataset comprises $5846$ measurements, of which $1755$ measurements are observational, and $4091$ measurements are from the five different single node interventions. We here compare the predicted order of Intersort, using the Wasserstein distance and an epsilon value of $1.5$, to the consensus true network consisting of $17$ edges (Sachs et al., 2005). Intersort achieves a $D_{top}$ of $3$, where we color the $3$ edges in the wrong direction in red.
@@ Appendix G Computational resources
The simulations of fig. 1 were run on an Apple M1 Macbook pro. The experiments for figs. 2 and 7 were run on a cluster with 20 CPUs, 16Gb of memory per CPU.